# How to Train Neural Field Representations A Comprehensive Study and Benchmark

### v1

@@ Abstract
In this work, we propose fit-a-nef [^1] GitHub repository available: https://github.com/samuelepapa/fit-a-nef, a JAX-based library that leverages parallelization to enable fast optimization of large-scale NeF datasets, resulting in a significant speed-up. With this library, we perform a comprehensive study that investigates the effects of different hyperparameters –including initialization, network architecture, and optimization strategies– on fitting NeFs for downstream tasks. Our study provides valuable insights on how to train NeFs and offers guidance for optimizing their effectiveness in downstream applications. Finally, based on the proposed library and our analysis, we propose Neural Field Arena, a benchmark consisting of neural field variants of popular vision datasets, including MNIST, CIFAR, variants of ImageNet, and ShapeNetv2. Our library and the Neural Field Arena will be open-sourced to introduce standardized benchmarking and promote further research on neural fields.
Figure 1: To investigate the use of Neural Fields (NeFs) as representations, we propose Fit-a-NeF (0(a)), a JAX-based library for efficient parallelized fitting of NeFs to datasets of signals, obtaining a $00\times1$ , 300 × speed-up. Owing to this efficiency, we are able to uncover the impact of NeF hyperparameter choices on their usability as representations – evaluated in downstream classification (Fig. 0(b)) – obtaining two important insights. First, 0(b)-left, it is vital to group NeFs in parameter-space, which we propose to enforce by sharing their network initializations. Second, 0(b)-right, improved reconstruction quality does not necessarily result in improved representation quality, implying an optimal combination of NeF expressivity and optimization for learning on NeFs. Incorporating these insights, we create a suite of NeF-based variants of classical CV datasets, Fig. 0(c). We bundle these Neural Datasets into a benchmark for learning on Neural Fields – which we name Neural Field Arena – hoping to enable standardized comparison in order to promote further research into this field.
@@ 1 Introduction
Neural Fields (NeFs) [46] have recently emerged as a versatile method for reconstructing discretely sampled continuous signals, with great success on various modalities from images [43, 42, 44], videos [35], 3D scenes and shapes [30, 34, 31, 49], to medical imaging [51, 33], audio [42, 18], and physical systems [39, 48, 23]. As coordinate-based models $f_{\theta}\colon x\mapsto y,x\in\mathbb{R}^{d},y\in\mathbb{R}^{c}$ (e.g. $d=2$ for images and $d=3$ for 3D shapes), their continuous nature allows for modeling signals without the limitations in resolution inherent in discrete representations; their capacity for modeling fine detail is only limited by the expressivity of the underlying NeF architecture. Furthermore, some modalities, such as scenes and shapes, do not have array representations that can be easily adapted to existing deep learning methods. Doubling down on their flexibility, a growing body of literature [14, 3, 32, 52, 16] has further proposed to use NeF parameterizations $f_{\theta}$ directly as representations of data. They are used to model images, videos, or shapes not as tensors but as functions on which to run downstream tasks (e.g., classify the function of an image instead of the pixel tensor of an image). In this setting, the parameters $\theta$ of a NeF – sometimes along with their computational structure – serve as input for a downstream model.
Despite their promise, NeFs lag behind as representations. Prior work has shown that performance on well-known benchmarks such as MNIST and CIFAR10 is far from state-of-the-art [32, 52]. We argue that two challenges contribute to their lack of performance: a practical one, and a theoretical one. The practical challenge is that a NeF is a neural network fit on a single discretely sampled signal, for instance, the pixels of an image. Fitting a new NeF for every sample in a dataset means training a new neural network from scratch, thus scaling up to whole datasets – especially large datasets with high-resolution signals – is prohibitively time-consuming [14, 32, 16]. The theoretical challenge stems from the fact that in this setting we treat the parameters of a neural network as a signal representation, and given our lack of understanding on how neural networks behave as a function of their architecture and optimization, we do not know what makes the parameters of a NeF a good representation for downstream tasks. Moreover, the scaling inefficiencies described above hinder us from forming hypotheses and testing them experimentally, inhibiting systematic research on NeF representations.
In this work, we address both challenges. Firstly, the existing libraries for learning neural networks are not necessarily optimal for fitting NeFs. Instead, we want a library that parallelizes the fitting of thousands of NeFs. We propose a JAX-based library [6, 20] - which we dub Fit-a-NeF - that leverages both in-device and across-device parallelization and reduces the time required for NeF fitting drastically.
@@
Following our previous argument, we fit NeFs on popular computer vision datasets, which we call Neural Datasets, starting with MNIST, CIFAR, and a modified version of ImageNet, and ShapeNet which we will continuously expand. To foster further research on NeF representations, we set up Neural Field Arena, a series of benchmarks on these Neural Datasets analogous to the benchmarks for the original datasets. In short, our contributions are as follows.
  • We provide a JAX-based library to quickly and easily fit millions of neural fields, for different modalities, including images and shapes, using accelerators with parallelization on and across several devices.
@@ 2 Representations in Parameter Space
To be more specific, for any two functions $f_{1}$ and $f_{2}$ (which we would like to fit with NeFs) that are $\epsilon-$close in function space, it is not always theoretically possible to fit neural networks to these respective functions that are also close in their parameter space [37]. In other words, two signals that are similar might not have NeF representations which are themselves similar. Furthermore, the same neural network can be realized with parameters that are far apart in parameter space [19, 7, 1]. Importantly, a recent line of work shows that it is possible to align such neural networks that are close in function space but distant in parameter space, for example by finding optimal permutations in the weight space [1] or finding optimal transport solutions to (soft) neuron alignment [41].
Together, these properties dramatically complicate the task of learning on NeFs. In the next section, we form hypotheses based on these properties for improving NeF parameterizations as downstream representations.
@@ 2.1 Grouping Representations in Parameter Space
Ainsworth et al. [1] posit that –modulo the permutation symmetries naturally present in neural networks– the local minima reached by optimization of different randomly initialized neural network weights form a convex basin. Following these insights, [53, 32, 52] show that downstream models that are invariant to these ambiguities improve performance dramatically over models that do not account for them. Both of these findings support the intuition that enforcing alignment in parameter space simplifies downstream learning. Our hypothesis is the smaller pairwise distances in parameter space may simplify training because grouped NeF representations span a smaller region in parameter space, preventing, at least partially, very different NeFs from corresponding to semantically similar signals.
@@
RQ1: Does sharing a random initialization help with learning better NeF representations?
Neural Dataset
Reconstruction Quality
Low
Medium
High
Ground Truth
Neural MNIST
Neural CIFAR10
Neural MicroImageNet
Neural ShapeNet
@@ 2.2.1 Overtraining Neural Fields
RQ2: Can overtraining NeFs negatively affect representation quality even if it improves reconstruction quality?
@@ 3.1 Fit-a-NeF: Fitting neural fields at scale
Speed comparisons. We show comparisons of Fit-a-NeF with a naive approach in Fig. 3. Specifically, we compare the time it takes to fit various neural datasets with our library and with a naive approach consisting of fitting one NeF at a time, on the same GPU. The gain depends on various factors such as the GPU memory and the size of the datapoints. For example, we can fit Neural MNIST in 17 seconds222this is true if the GPU memory allows for training of the whole MNIST in parallel, e.g., on an A100 vs. the 3 hours it would take with the naive approach. In short, with Fit-a-NeF it took 62 GPU days on A100s to run all the experiments reported in this work, whilst a naive approach would have required approximately 30 GPU years.
@@ 3.2 Datasets and Metrics
In order to address our research questions and establish the quality of NeF representations, we train a downstream model for downstream tasks. More specifically, we choose NeF classification as a downstream task and report test accuracy, following prior work on representation learning [4, 10, 27]. For a downstream model, it is important to use a method that respects the neural network permutation symmetries. We choose the work of Zhang et al. [52] that represents neural fields as computational graphs and processes them with standard graph networks. We create a variant that relies on a simpler message passing neural network (MPNN). We provide details about the classifier implementation in the appendix.
@@
For the quality of reconstruction or generation, the metric of choice is the peak signal-to-noise ratio (PSNR) in images and intersection over union (IOU) in shapes.
We propose to extend both to what we call their off-grid counterparts; they are calculated using coordinates that are different from those used to fit the NeF to the signal. For images, the off-grids points lie in-between pixels, and the ground-truth image is obtained through a $\beta^{(1)}$-spline interpolation of the original image at those coordinates. For shapes, new points and their respective occupancy is calculated for all shapes in the dataset.
We refer to these metrics as “off-grid-PSNR” and “off-grid-IOU”, and use PSNR and IOU only to refer to their on-grid versions, which are computed on the coordinates used to fit the signals. In the experiments, we find that the ratio of the two, that is $\frac{\texttt{off-grid-PSNR}}{\texttt{on-grid-PSNR}}$, and $\frac{\texttt{off-grid-IOU}}{\texttt{on-grid-IOU}}$, is particularly informative.
@@ 3.4 Architecture
In the appendix, we additionally experiment with different architectures classically used in NeF literature, SIREN [42], RFFNets [44], and MFNs (FourierNet) [17] to assess their impact on representation quality, as well as the effects of the learning rate and weight decay. Given a certain architecture, we also explore its depth and width.
Figure 4: To make this plot we have fitted 10 ShapeNet-10 Neural Datasets varying initialization and total number of steps. The histograms show the distribution of pairwise distances of the NeF representations. It is evident that shared initialization produces more grouped representations, and that the pairwise distances increase with the number of steps.
@@ 4 Results and discussion
4.1 Experiment 1: Shared vs. Random Initialization
@@
Figure 7: We fit 220 neural datasets using different hidden dimensions and the number of steps while keeping the same shared initialization. We find that the ratio of off-grid reconstruction quality and in-grid reconstruction quality can be used to form a heuristic that correlates with high test accuracy
4.2 Experiment 2: Overtraining Neural Fields
@@
The PSNR and IOU measure the reconstruction quality of neural fields, whereas the off-grid PSNR and off-grid IOU measure overfitting. We plot the classification accuracy achieved with neural datasets of different reconstruction quality in Fig. 6. We observe that the classification accuracy increases with reconstruction quality but then reaches a maximum and starts decreasing. The general trend is consistent across different datasets, architectures, and optimization choices. In Fig. 7 we plot the classification accuracy against the ratio $\frac{\texttt{off-grid-PSNR}}{\texttt{on-grid-PSNR}}$ for images or $\frac{\texttt{off-grid-IOU}}{\texttt{on-grid-IOU}}$ for shapes. Interestingly, a clear pattern emerges that is consistent across datasets: the quality of the representations keeps increasing with the number of steps until the ratio starts decreasing. We conclude that overfitting negatively affects downstream performances.
4.3 Experiment 3: Expressivity of Neural Field Architectures
@@
Figure 8 aggregates results from 110 experiments across the different datasets and shows how increasing the hidden dimension tends to lead to worse representation quality while the higher expressivity of the NeF leads to higher PSNR.
@@ 5 Benchmarks
Part of the fast progress of Deep Learning is due to publicly available datasets for benchmarking that allow for fair comparisons. Neural representations are a relatively new approach to representation learning and because of that, we lack good benchmark datasets to compare architectures that perform different downstream tasks on them. Since we observed great variibility in downstream performance as we vary the NeF architecture and optimization strategy, proper benchmarking is key to systematic progress hereupon.
After having explored the best optimization and architecture for several datasets, we now establish a new benchmark with existing downstream models that had previously achieved the state-of-the-art in classification using neural representations.
@@
Table 3: We benchmark the new three proposed neural datasets using state-of-the-art methods from literature [32, 52]. We obtain
@@ 6 Related work
Recent years have seen a surging interest in the development of methods to facilitate NeF-based signal reconstructions. A number of works explore changes in neural network architectures to address the spectral bias of conventional ReLU MLPs towards low-frequency signals [38], which impedes the learning of high-fidelity signal reconstructions in fine detail. Tancik et al. [44] use a random Fourier feature mapping applied to the input of a NeF to prevent over smoothing, whereas Sitzmann et al. [42] propose to use periodic activations to artificially inflate the frequency spectrum of the network, achieving better modeling of fine details. Multiplicative filters networks [17] are a different class of functions that avoids traditional compositional network depth and multiplies together simple functions, which has the added benefit of direct control over the network’s frequency spectrum [40]. Importantly, all of these works only investigate NeF architectures for improved signal reconstruction quality. We instead focus on researching NeF architectures for their viability as downstream representations, i.e. learning on the parameter space of NeFs.
A different direction of research attempts to alleviate the computational burden of fitting individual NeFs to dataset samples, and to reduce the complexity of a NeF as representation. Building on the concept of neural network conditioning through modulation [36], Chan et al. [8], Mehta et al. [29] propose to use a single NeF to represent a diversity of signals. The activations of this NeF, which shares parameters across the entire dataset, are modulated with datapoint-specific low-dimensional latents to yield datapoint-specific outputs. The latents may be obtained using an encoder [29], or learned by using a combination of meta-learning and auto-decoding [34]. Dupont et al. [14] provided the insight that these learned latents may be used as downstream representation instead of the parameters of a sample-specific NeF to simplify the task of learning on a parameter space. Although reducing the complexity of learning on parameter space, this approach limits performance – both in terms of signal reconstruction quality and downstream performance [3]. Bauer et al. [3] show that a spatial grid of latent vectors is needed to reach competitive downstream performance, entailing all limitations that come with discrete data representations. In this work – to obtain the most general and flexible signal parameterizations – we instead focused on NeFs without shared parameters or sample-specific conditioning and explore what factors facilitate learning on their parameter space directly in downstream tasks.
@@
We briefly mention a few applications of neural representations. Several works propose the use of neural representations for generating shapes using diffusion models [16]. In the context of dynamical systems and PDEs, [48] show that neural representations can be used effectively for dynamics forecasting to the flow of partial differential equations as spatially and time-continuous functions.
@@
The exploration of deep learning architectures for processing neural network parameters is relatively recent. Here, we offer an overview of the key studies in this domain. The studies by [15] and [45] revolve around predicting characteristics of trained neural networks (NNs) by scrutinizing their weight configurations. [15] is primarily focused on estimating the hyperparameters used during network training, while [45] is dedicated to evaluating a network’s generalization capabilities. Both investigations involve applying conventional neural networks to either the flattened weight data or their statistical attributes. On the other hand, [47] introduced a novel concept where NNs are processed using another neural network that operates on a combination of their high-order spatial derivatives. This technique is particularly suitable for NeF representations where derivative information plays a crucial role. Nevertheless, the adaptability of these networks to broader tasks remains uncertain, and the demand for high-order derivatives can introduce significant computational overhead. Lastly, the works of [52], proposed a transformer-like architecture for compatibility with multiplicative filters networks, and the studies by [32] and [53], which initially emphasized the importance of augmentations and permutation-invariant architectures for processing neural field weights.

### v2

# How to Train Neural Field Representations: A Comprehensive Study and Benchmark

@@ Abstract
In this work, we propose a JAX-based library [^1] https://fit-a-nef.github.io/ that leverages parallelization to enable fast optimization of large-scale NeF datasets, resulting in a significant speed-up. With this library, we perform a comprehensive study that investigates the effects of different hyperparameters on fitting NeFs for downstream tasks. In particular, we explore the use of a shared initialization, the effects of overtraining, and the expressiveness of the network architectures used. Our study provides valuable insights on how to train NeFs and offers guidance for optimizing their effectiveness in downstream applications. Finally, based on the proposed library and our analysis, we propose Neural Field Arena, a benchmark consisting of neural field variants of popular vision datasets, including MNIST, CIFAR, variants of ImageNet, and ShapeNetv2. Our library and the Neural Field Arena will be open-sourced to introduce standardized benchmarking and promote further research on neural fields.
Figure 1: To investigate the use of Neural Fields (NeFs) as representations, we propose Fit-a-NeF (1(a)), a JAX-based library for efficient parallelized fitting of NeFs to datasets of signals, obtaining a $00\times1$ , 300 × speed-up. Owing to this efficiency, we are able to uncover the impact of NeF hyperparameter choices on their usability as representations – evaluated in downstream classification (Fig. 1(b)) – obtaining two important insights. First, 1(b)-left, it is vital to group NeFs in parameter-space, which we propose to enforce by sharing their network initializations. Second, 1(b)-right, improved reconstruction quality does not necessarily result in improved representation quality, implying an optimal combination of NeF expressivity and optimization for learning on NeFs. Incorporating these insights, we create a suite of NeF-based variants of classical CV datasets, Fig. 1(c). We bundle these Neural Datasets into a benchmark for learning on Neural Fields – which we name Neural Field Arena – hoping to enable standardized comparison in order to promote further research into this field.
@@ 1 Introduction
Neural Fields (NeFs) [48] have recently emerged as a versatile method for reconstructing discretely sampled continuous signals, with great success on various modalities from images [45, 44, 46], videos [37], 3D scenes and shapes [32, 36, 33, 51], to medical imaging [53, 35], audio [44, 18], and physical systems [41, 50, 24]. The continuous nature of coordinate-based models $f_{\theta}\colon x\mapsto y,x\in\mathbb{R}^{d},y\in\mathbb{R}^{c}$ (e.g. $d=2$ for images and $d=3$ for 3D shapes), allows for modeling signals without the limitations in resolution inherent in discrete representations; their capacity for modeling fine detail is only limited by the expressivity of the underlying NeF architecture. Furthermore, some modalities, such as scenes and shapes, do not have array representations that can be easily adapted to existing deep learning methods. Doubling down on their flexibility, a growing body of literature [14, 3, 34, 25, 16] has further proposed to use NeF parameterizations $f_{\theta}$ directly as representations of data. They are used to model images, videos, or shapes not as tensors but as functions on which to run downstream tasks (e.g., classify the function of an image instead of the pixel tensor of an image). In this setting, the parameters $\theta$ of a NeF –sometimes along with their computational structure– serve as input for a downstream model.
Despite their promise, NeFs lag behind as image representations, while showing promise for shape datasets [34, 25]. We argue that two challenges contribute to their lack of performance: a practical one, and a theoretical one. The practical challenge is that a NeF is a neural network fit on a single discretely sampled signal, for instance, the pixels of an image. Fitting a new NeF for every sample in a dataset means training a new neural network from scratch, thus scaling up to whole datasets –especially large datasets with high-resolution signals– is prohibitively time-consuming [14, 34, 16]. The theoretical challenge stems from the fact that in this setting we treat the parameters of a neural network as a signal representation, and given our lack of understanding of how neural networks behave as a function of their architecture and optimization, we do not know what makes the parameters of a NeF a good representation for downstream tasks. Moreover, the scaling inefficiencies described above hinder us from forming hypotheses and testing them experimentally, inhibiting systematic research on NeF representations.
In this work, we address both challenges. Firstly, the existing libraries for learning neural networks are not necessarily optimal for fitting NeFs. To this end, we propose a JAX-based library [6, 20] –which we dub Fit-a-NeF– that leverages both in-device and across-device parallelization to reduce the time required for NeF fitting drastically.
@@
Following our previous argument, we fit NeFs on popular datasets, which we call Neural Datasets, starting with MNIST, CIFAR, a modified version of ImageNet, and ShapeNet. To foster further research on NeF representations, we set up Neural Field Arena, a series of benchmarks on these Neural Datasets analogous to the benchmarks for the original datasets. In short, our contributions are as follows.
  • We provide a JAX-based library to quickly and easily fit millions of neural fields for different modalities, including images and shapes, using accelerators with parallelization on and across several devices.
@@ 2 Representations in Parameter Space
To be more specific, for any two functions $f_{1}$ and $f_{2}$ (which we would like to fit with NeFs) that are $\epsilon-$close in function space, it is not always theoretically possible to fit neural networks to these respective functions that are also close in their parameter space [39]. In other words, two signals that are similar might not have NeF representations which are themselves similar. Furthermore, the same neural network can be realized with parameters that are far apart in parameter space [19, 7, 1]. Importantly, a recent line of work shows that it is possible to align such neural networks that are close in function space but distant in parameter space, for example by finding optimal permutations in the weight space [1] or finding optimal transport solutions to (soft) neuron alignment [43]. Together, these properties dramatically complicate the task of learning on NeFs.
@@ 2.1 Grouping Representations in Parameter Space
Ainsworth et al. [1] posit that –modulo the permutation symmetries naturally present in neural networks– the local minima reached by optimization of different randomly initialized neural network weights form a convex basin. Following these insights, [54, 34, 25] show that downstream models that are invariant to these ambiguities improve performance dramatically over models that do not account for them. Both of these findings support the intuition that enforcing alignment in parameter space simplifies downstream learning. Our hypothesis is that the smaller pairwise distances in parameter space may simplify training because grouped NeF representations span a smaller region in parameter space, preventing, at least partially, very different NeFs from corresponding to semantically similar signals.
@@
RQ1: Do neural datasets obtained starting from a shared initialization lead to better downstream performance, when measured in terms of accuracy?
@@ 2.2.1 Overtraining Neural Fields
RQ2: Can overtraining NeFs negatively affect representation quality in terms of downstream accuracy, even if it improves the reconstruction of the given signals?
Neural Dataset Reconstruction Quality
Low Medium High Ground Truth
Neural MNIST Refer to caption Refer to caption Refer to caption Refer to caption
Neural CIFAR10 Refer to caption Refer to caption Refer to caption Refer to caption
Neural MicroImageNet Refer to caption Refer to caption Refer to caption Refer to caption
Neural ShapeNet Refer to caption Refer to caption Refer to caption Refer to caption
@@ 3.1 Fit-a-NeF: Fitting Neural Fields at Scale
Speed comparisons. We show comparisons of Fit-a-NeF with a naive approach in Fig. 3. Specifically, we compare the time it takes to fit various neural datasets with our library and with a naive approach consisting of fitting one NeF at a time, on the same GPU. The gain depends on various factors such as the GPU memory and the size of the datapoints. For example, we can fit Neural MNIST in 17 seconds222this is true if the GPU memory allows you to train the whole MNIST in parallel, e.g., on an NVIDIA A100 vs. the 3 hours it would take with the naive approach. In short, with Fit-a-NeF it took 62 GPU days on A100s to run all the experiments reported in this work, whilst a naive approach would have required approximately 30 GPU years.
@@ 3.2 Datasets and Metrics
In order to address our research questions and establish the quality of NeF representations, we train a downstream model for downstream tasks. More specifically, we choose NeF classification as a downstream task and report test accuracy, following prior work on representation learning [4, 10, 29]. For a downstream model, it is important to use a method that respects the neural network permutation symmetries. We choose the work of Kofinas et al. [25] that represents neural fields as computational graphs and processes them with standard graph networks. We create a variant that relies on a simpler message passing neural network (MPNN). We provide details about the classifier implementation in the appendix.
@@
We use peak signal-to-noise ratio (PSNR) in images and intersection over union (IOU) in shapes to measure the quality of reconstruction or generation. Additionally, we propose to extend both to what we call their off-grid counterparts; calculated using coordinates that are different from those used to fit the NeF to the signal. For images, the off-grids points lie in-between pixels, and the ground-truth image is obtained through a $\beta^{(1)}$-spline interpolation of the original image at those coordinates. For shapes, new points and their respective occupancy is calculated for all shapes in the dataset.
We refer to them as “off-grid-PSNR” and “off-grid-IOU”, and use PSNR and IOU only to refer to their on-grid versions, which are computed on the coordinates used to fit the signals. In the experiments, we find the ratio between the two metrics $\frac{\texttt{off-grid-PSNR}}{\texttt{on-grid-PSNR}}$, and $\frac{\texttt{off-grid-IOU}}{\texttt{on-grid-IOU}}$ particularly informative.
@@ 3.4 Architecture
In the main text of the paper we using SIREN [44]. In the supplementary material, we additionally experiment with different architectures classically used in NeF literature, RFFNets [46], and MFNs (FourierNet) [17] to assess their impact on representation quality, as well as the effects of the learning rate and weight decay. Given a certain architecture, we also explore its depth and width.
Figure 4: The histograms show the distribution of pairwise distances of the NeF representations. We fit 10 ShapeNet-10 Neural Datasets varying initialization and total number of steps. Shared initialization produces more grouped representations, and pairwise distances increase with the number of steps.
@@ 4 Results and Discussion
4.1 Shared vs. Random Initialization
@@
Figure 7: We fit 220 Neural Datasets using different hidden dimensions and the number of steps while keeping the same shared initialization. We find that the ratio of off-grid reconstruction quality and in-grid reconstruction quality can be used to form a heuristic that correlates with high test accuracy.
4.2 Overtraining Neural Fields
@@
The PSNR and IOU measure the reconstruction quality of neural fields, whereas the off-grid PSNR and off-grid IOU measure overfitting. We plot the classification accuracy achieved with Neural Datasets of different reconstruction quality in Fig. 6. We observe that the classification accuracy increases with reconstruction quality but then reaches a maximum and starts decreasing. The general trend is consistent across different datasets, architectures, and optimization choices. In Fig. 7 we plot the classification accuracy against the ratio $\frac{\texttt{off-grid-PSNR}}{\texttt{on-grid-PSNR}}$ for images or $\frac{\texttt{off-grid-IOU}}{\texttt{on-grid-IOU}}$ for shapes. Interestingly, a clear pattern emerges that is consistent across datasets: when the number of steps is the only hyperparameter that is changed, stopping as soon as the NeF starts showing signs of overtraining leads to the best representation quality. We conclude that overfitting negatively affects downstream performances.
4.3 Expressivity of Neural Field Architectures
@@
Figure 8 aggregates results from 110 experiments across the different datasets and shows how increasing the hidden dimension tends to lead to worse representation quality while the higher expressivity of the NeF leads to higher PSNR. When exploring different activation functions and model layouts (see supplementary material) we observe a consistent behavior.
@@ 5 Benchmarks
Part of the fast progress of Deep Learning is due to publicly available datasets for benchmarking that allow for fair comparisons. Neural representations are a relatively new approach to representation learning and because of that, we lack good benchmark datasets to compare architectures that perform different downstream tasks on them. Since we observed great variability in downstream performance as we vary the NeF architecture and optimization strategy, proper benchmarking is key to systematic progress hereupon.
After having explored the best optimization and architecture for several datasets, we now establish a new benchmark with existing downstream models that had previously achieved the state-of-the-art classification using neural representations.
@@
Table 3: We benchmark the new three proposed neural datasets using state-of-the-art methods from literature [34, 25].
@@ 6 Related Work
Recent years have seen a surging interest in the development of methods to facilitate NeF-based signal reconstructions. A number of works explore changes in neural network architectures to address the spectral bias of conventional ReLU MLPs towards low-frequency signals [40], which impedes the learning of high-fidelity signal reconstructions in fine detail. Tancik et al. [46] use a random Fourier feature mapping applied to the input of a NeF to prevent over-smoothing, whereas Sitzmann et al. [44] propose to use periodic activations to artificially inflate the frequency spectrum of the network, achieving better modeling of fine details. Multiplicative filters networks [17] are a different class of functions that avoids traditional compositional network depth and multiplies together simple functions, which has the added benefit of direct control over the network’s frequency spectrum [42]. Importantly, all of these works only investigate NeF architectures for improved signal reconstruction quality. We instead focus on researching NeF architectures for their viability as downstream representations, i.e. learning on the parameter space of NeFs.
A different direction of research attempts to alleviate the computational burden of fitting individual NeFs to dataset samples, and to reduce the complexity of a NeF as representation. Building on the concept of neural network conditioning through modulation [38], Chan et al. [8], Mehta et al. [31] propose to use a single NeF to represent a diversity of signals. The activations of this NeF, which shares parameters across the entire dataset, are modulated with datapoint-specific low-dimensional latents to yield datapoint-specific outputs. The latents may be obtained using an encoder [31], or learned by using a combination of meta-learning and auto-decoding [36]. Dupont et al. [14] provided the insight that these learned latents may be used as downstream representation instead of the parameters of a sample-specific NeF to simplify the task of learning on a parameter space. Although reducing the complexity of learning on parameter space, this approach limits performance – both in terms of signal reconstruction quality and downstream performance [3]. Bauer et al. [3] show that a spatial grid of latent vectors is needed to reach competitive downstream performance, entailing all limitations that come with discrete data representations. In this work –to obtain the most general and flexible signal parameterizations– we instead focused on NeFs without shared parameters or sample-specific conditioning and explored what factors facilitate learning on their parameter space directly in downstream tasks.
@@
We briefly mention a few applications of neural representations. Several works propose their use for generating shapes using diffusion models [16]. In the context of dynamical systems and PDEs, [50] show that neural representations can be used for dynamics forecasting to the flow of partial differential equations as spatially and time-continuous functions.
@@
The exploration of deep learning architectures for processing neural network parameters is relatively recent. Here, we offer an overview of the key studies. The works by [15] and [47] revolve around predicting characteristics of trained neural networks (NNs) by scrutinizing their weight configurations. [15] is primarily focused on estimating the hyperparameters used during network training, while [47] is dedicated to evaluating a network’s generalization capabilities. Both investigations involve applying conventional neural networks to either the flattened weight data or their statistical attributes. On the other hand, [49] introduced a novel concept where NNs are processed using another neural network that operates on a combination of their high-order spatial derivatives. This technique is particularly suitable for NeF representations where derivative information plays a crucial role. Nevertheless, the adaptability of these networks to broader tasks remains uncertain, and the demand for high-order derivatives can introduce significant computational overhead. Lastly, the works of [25], proposed a transformer-like architecture for compatibility with multiplicative filters networks, and the studies by [34] and [54], which initially emphasized the importance of augmentations and permutation-invariant architectures for processing neural field weights.